[Hallc_running] RC Daily Update
tue81317 at temple.edu
Wed Feb 27 15:12:58 EST 2019
Running has been fairly smooth since beam was restored around 7pm, but with
limited current. For most of the night we could only run at 30uA to
40uA. Recently we have come up to 50uA and we're pushing to get a little
more. Some magnet issues overnight delayed us for an hour or so. We are
running Phase-2 kinematics now and through tomorrow.
For more information, the RC meeting will be held today at 4:30 in the
counting house. For
updates and meeting details see:
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running